# Spectral Properties of Irradiated Circumbinary Disks around Binary Black Holes Governed by Hydrogen Opacities Dependent on Temperature and Density

### v1

@@ 3.2 Energy equation
Next, the irradiation heating rate of the CBD is given by equation (16) of Lee et al. (2024) as
$Q_{\rm irr}=\frac{A_{1}L_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\Biggr{[}(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\Biggr{]},\$ (41)
where $A_{1}$ and $A_{2}$ are the absorption fractions of the incident radiation energy from the CPD and the CSD, respectively, $Q_{12}=A_{2}L_{2}/(A_{1}L_{1})$, and $\beta_{1}$ and $\beta_{2}$ are the parameters given by
$\beta_{1}=\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{q^{2}}{(1+q)^{2}},\,\,\beta_{2}=\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{1}{(1+q)^{2}},$ (42)
respectively. Note that $\beta_{1}$ and $\beta_{2}$ are less than unity for $q\leq 1$. Since $A_{1}$ and $A_{2}$ are unlikely to differ significantly for two minidisks in similar physical states, we assume that $A_{1}=A_{2}=A$, resulting in $Q_{12}=q$.
@@
$\frac{3}{4\pi}\frac{GM\dot{M}}{r_{\rm in}^{3}}\frac{1}{\xi^{3}}+\frac{AL_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\left[(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\right]=\frac{32}{3}\frac{1}{(1+x_{e})^{4}}\frac{\sigma{T_{\rm 0}^{4}}}{\tau}\frac{Y^{8}}{\xi^{4}}.$ (43)
Model Opacity ($\kappa$) Optical depth ($\tau$) Temperature relation Outer boundary condition ($Y_{\rm out}$) Remarks
Model 0 -- $16/3$ $T_{\rm s}=T_{\rm c}$ Equation (29) of Lee et al. (2024) Lee et al. (2024)
Model 1 $\kappa_{\rm es}$ Equation (36) $T_{\rm s}\neq T_{\rm c}$ Equation (50) --
Model 2 $\sqrt{\kappa_{\rm es}\kappa_{\rm ff}}$ Equation (37) $T_{\rm s}\neq T_{\rm c}$ Equation (51) --
Model 3 $\sqrt{\kappa_{\rm a}\left(\kappa_{\rm a}+\kappa_{\rm es}\right)}$ Equation (38) $T_{\rm s}\neq T_{\rm c}$ Equation (LABEL:eq:yout3) --
Table 1: Summary of the models incorporating three different opacity prescriptions. The columns, from left to right, list the model identifier, the adopted opacity $\kappa$, the expression for the optical depth $\tau$, the relation between the surface temperature $T_{\rm s}$ and midplane temperature $T_{\rm c}$, the outer boundary condition $Y_{\rm out}$, and additional remarks. Model 0 reproduces the fiducial case from Lee et al. (2024), assuming $T_{\rm s}=T_{\rm c}$ without employing any physical opacity. Model 1 adopts pure electron scattering opacity. Model 2 describes the regime where electron scattering dominates over free-free absorption. Model 3 includes an effective opacity accounting for both free-free and bound-free absorption and electron-scattering.
3.3 Basic equation
@@
Our model has seven parameters: $M$, $\dot{m}$, $a_{0}$, $C_{\rm gap}$, $q$, $A$, and $\xi_{\rm out}$. Among them, the following five parameters are fixed throughout the paper as $\dot{m}=1$, $a_{0}=1000$, $C_{\rm gap}=2$, $q=1$, and $A=0.1$. The dependencies of the CBD temperature and spectra on the remaining two parameters are examined in the next section.
4 Results
@@
4.1 Circumbinary Disk Temperature Profiles
@@
4.2 Circumbinary Disk Spectra
@@
[…]
where $\delta$ is the inclination angle of the disk and $D$ is the distance between the source and the earth. In the following we adopt $\delta=0$ unless otherwise noted.
Figure 3: SEDs of the CBD for the four opacity models listed in Table 1. Panels (a)–(c) correspond to total binary masses of $100\,M_{\odot}$, $10\,M_{\odot}$ (fiducial), and $1000\,M_{\odot}$, respectively. Each panel plots the normalized luminosity $\nu L_{\nu}/L_{\rm Edd}$ versus frequency on logarithmic axes. Solid curves indicate Model 0 (black), Model 1 (blue), Model 2 (green), and Model 3 (red); the dashed black curve reproduces Model 0 without external irradiation.
Figure 4: Same format as Figure 3 but for different outer disk radii, varying the outer radius $\xi_{\rm out}$ while keeping the binary mass fixed at $M=100\,M_{\odot}$. Panels (a)–(c) adopt $\xi_{\rm out}=10^{3}$, $10^{4}$, and $10^{5}$, respectively; all other parameters are identical to those in Figure 3.
@@
Varying the total binary mass primarily shifts the entire SED along the frequency axis. For a lower-mass system ($M=10\,M_{\odot}$; Panel b), both the irradiation-induced IR peak and the viscously heated ultraviolet (UV) shoulder shift to higher frequencies. In contrast, a higher-mass system ($M=1000\,M_{\odot}$; Panel c) produces a cooler and more extended CBD, moving both features to lower frequencies.
At high frequencies ($\nu\gtrsim 10^{15}\,\mathrm{Hz}$), all four models yield nearly identical spectra. In this regime, viscous heating dominates, and the surface temperature is governed by the local energy dissipation rate, making the spectral output largely independent of the assumed opacity.
At lower frequencies, the IR peak shows a non-monotonic trend in flux across the models. Model 0, which assumes $T_{\mathrm{s}}=T_{\mathrm{c}}$ and ignores all physical opacity sources, produces the brightest spectrum. This is because the surface temperature is set equal to the higher midplane temperature, and radiation escapes without being absorbed. Among the models that include opacity, Model 3 gives the strongest IR flux, followed by Model 2 and then Model 1. This order reflects differences in surface temperature profiles (see Figure 2), which are caused by how efficiently each model absorbs and thermalizes the incoming radiation in the outer part of the disk.
@@
In Model 3, the inclusion of bound–free absorption greatly increases the Rosseland mean opacity in the temperature range $T\sim 10^{3}$$10^{4}\,\mathrm{K}$, which coincides with the irradiated outer disk regions where the disk flares. This enhances the local trapping of irradiating photons and raises the surface temperature, leading to stronger IR re-emission than in Model 2, where only free–free opacity is present. In contrast, electron scattering only in Model 1 does not effectively trap irradiating photons due to the weaker disk flaring, so the surface remains cooler and the IR flux is lower.
These trends indicate that observational constraints on CBD opacity sources are most effective in the IR for high-mass binaries and in the optical for low-mass binaries, where the spectral differences among models are most pronounced.
@@
In Panel (b) of Figure 3, the spectrum of Model 3 (red curve) clearly exceeds those of the other three models across the mid-IR through the optical and near-ultraviolet frequencies. Only in the far-IR (frequencies $\nu\lesssim 10^{13.5}\,\mathrm{Hz}$) does Model 0 (black curve), which includes no physical absorption opacity, slightly surpass Model 3. Thus, within a single panel, a two-tier behavior emerges: Model 0 leads only at the lowest frequencies, while Model 3 dominates elsewhere. Furthermore, when the SEDs of Model 3 are compared across the three mass cases (Panels a, b, and c), the $10\,M_{\odot}$ case exhibits the highest normalized luminosity.
This behavior arises from the combination of the CBD’s temperature profile and the reprocessing of irradiation from the two minidisks surrounding the binary. In the $10\,M_{\odot}$ system, the outer disk cools to temperatures below $10^{4}\,$, where hydrogen is partially ionized and bound–free opacity becomes especially strong. High-energy photons emitted by the two minidisks are absorbed in these outer layers, thermalized, and then re-emitted at IR wavelengths, producing a pronounced IR peak. Meanwhile, the inner disk remains hot enough that its surface temperature does not drop significantly, enhancing the mid- and high-frequency emission.
@@
The apparent “triple peaked” shape in the SED of Model 3 is in fact a modification of the basic double-peaked SED structure. The two primary peaks, one from viscous heating in the inner CBD and the other from reprocessed emission from the irradiated outer disk, are supplemented by a smaller subpeak in the mid-frequency range. This subpeak arises because bound–free absorption in Model 3 distorts the irradiated component of the spectrum, creating a visible peak between the main IR peak and the UV shoulder. Therefore, it is most appropriate to describe the SED as fundamentally double-peaked, with an additional mid-frequency peak caused by bound–free opacity effects.

### v2

@@ III.2 Energy equation
Next, the irradiation heating rate is determined by the irradiated flux on the surface of the CBD. The irradiated flux on the surface of the CBD is defined by equation (A12) of Lee et al. (2024) as
$F_{{\rm irr},i}=\frac{L_{i}}{4\pi d_{i}^{2}}(\vec{l}_{i}\cdot\vec{n})(\vec{n}\cdot\vec{e}_{z}),$ (42)
where $i=1$ and $i=2$ represent the primary and secondary black holes, respectively, $d_{i}$ is the distance between each irradiation source and the surface element on the CBD, $\vec{l}_{i}$ is the unit vector directed from the CBD surface toward the primary and secondary black holes, $\vec{n}$ denotes the unit vector normal to the CBD surface, and $\vec{e}_{z}$ is the unit vector along the $z$ axis (see also Appendix A.1 of Lee et al. (2024) for the detailed derivation of the irradiation flux).
The azimuthally averaged irradiation heating rate is then determined by (see also equation (16) of Lee et al. (2024))
$Q_{\rm irr}$ $=$ $2\int_{0}^{2\pi}\,A_{1}F_{\rm irr,1}+A_{2}F_{\rm irr,2}\,d\phi$
$=$ $\frac{A_{1}L_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\Biggr[(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\Biggr],$
where $A_{1}$ and $A_{2}$ are the absorption fractions of the incident radiation energy from the CPD and the CSD, respectively, $Q_{12}=A_{2}L_{2}/(A_{1}L_{1})$ with $L_{1}$ and $L_{2}$ being the bolometric luminosities defined in equation (16), and $\beta_{1}$ and $\beta_{2}$ are the parameters given by
$\beta_{1}$ $=$ $\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{q^{2}}{(1+q)^{2}},$ (44)
$\beta_{2}$ $=$ $\frac{3}{2}\frac{1}{C_{\rm gap}^{2}}\frac{1}{(1+q)^{2}},$ (45)
respectively. Note that $\beta_{1}$ and $\beta_{2}$ are less than unity for $q\leq 1$. Since $A_{1}$ and $A_{2}$ are unlikely to differ significantly for two minidisks in similar physical states, we assume that $A_{1}=A_{2}=A$, resulting in $Q_{12}=q$. Moreover, we treat $A$ as a constant. In the optically thick limit, the albedo 1A1-A is nearly independent of the local conditions and becomes constant, because multiple scattering and absorption events cause the reflected fraction to depend mainly on the ratio of scattering to absorption opacity rather than on detailed local variations (Fukue, 2018). As shown in equation (LABEL:eq:qirr), the irradiation heating rate depends not only on the radiated flux from the minidisks, represented by the term outside the brackets, but also on the flaring structure of the CBD, which appears through the terms $dY/d\xi$ and $Y/\xi$.
@@
$\frac{3}{4\pi}\frac{GM\dot{M}}{r_{\rm in}^{3}}\frac{1}{\xi^{3}}$
$+\frac{AL_{1}}{2\pi r_{\rm in}^{2}}\frac{1}{\xi}\left[(1+Q_{12})\frac{dY}{d\xi}-\frac{\beta_{1}+Q_{12}\beta_{2}}{\xi^{2}}\left(\frac{Y}{\xi}-\frac{1}{2}\frac{dY}{d\xi}\right)\right]$
$=\frac{32}{3}\frac{1}{(1+x_{e})^{4}}\frac{\sigma{T_{\rm 0}^{4}}}{\tau}\frac{Y^{8}}{\xi^{4}}.$ (46)
Although our equations remain 1D and axisymmetric, the binary imprint appears most clearly in the irradiation term $Q_{\rm irr}$. As defined by equation (LABEL:eq:qirr), $Q_{\rm irr}$ depends on the two minidisks’ luminosities $L_{1},L_{2}$, the absorption factors $A_{1},A_{2}$ (taken equal here), and the geometric coefficients $\beta_{1},\beta_{2}$, with the ratio $Q_{12}=A_{2}L_{2}/(A_{1}L_{1})$. The geometry of incidence, captured by the dot products in $F_{\rm{irr},\it{i}}$, together with $L_{1,2}$, $A_{1,2}$, and $\beta_{1,2}$, determines the intensity and spectral shape of the irradiation on the CBD surface. Therefore, the mass ratio $q$ and the partitioned accretion rates influence $Q_{\rm irr}$ through $L_{1,2}$, which in turn sets the outer–disk surface–temperature profile and the IR–optical SED. In our model, the CBD responds locally by satisfying the energy balance shown in equation (46).
Opacity model Opacity ($\kappa$) Optical depth ($\tau$) Temperature relation Outer boundary condition ($Y_{\rm out}$)
Model 0 - $16/3$ $T_{\rm s}=T_{\rm c}$ Equation (29) of Lee et al. (2024)
Model 1 $\kappa_{\rm es}$ Equation (37) $T_{\rm s}\neq T_{\rm c}$ Equation (53)
Model 2 $\sqrt{\kappa_{\rm es}\kappa_{\rm ff}}$ Equation (38) $T_{\rm s}\neq T_{\rm c}$ Equation (54)
Model 3 $\sqrt{\kappa_{\rm a}\left(\kappa_{\rm a}+\kappa_{\rm es}\right)}$ Equation (39) $T_{\rm s}\neq T_{\rm c}$ Equation (LABEL:eq:yout3)
Table 1: Summary of the models incorporating three different opacity prescriptions. The columns, from left to right, list the model identifier, the adopted opacity $\kappa$, the expression for the optical depth $\tau$, the relation between the surface temperature $T_{\rm s}$ and midplane temperature $T_{\rm c}$, the outer boundary condition $Y_{\rm out}$. Model 0 reproduces the fiducial case from Lee et al. (2024), assuming $T_{\rm s}=T_{\rm c}$ without employing any physical opacity. Model 1 adopts pure electron scattering opacity. Model 2 describes the regime where electron scattering dominates over free-free absorption. Model 3 includes an effective opacity accounting for both free-free and bound-free absorption and electron-scattering. For all models, only gas pressure is considered; radiation and magnetic pressures are neglected.
III.3 Basic equation
@@
III.4 Parameter dependence of CBD spectral shape
Our model has eight parameters: $q$, $\alpha_{\rm SS}$, $A$, $C_{\rm gap}$, $\dot{m}$, $M$, $a_{0}$, and $\xi_{\rm out}$. Throughout this paper, the fiducial values of these parameters are set as follows: $q=1$, $\alpha_{\rm SS}=0.1$, $C_{\rm gap}=2$, $A=0.1$, $\dot{m}=1$, $M=100\,M_{\odot}$, $a_{0}=1000$, and $\xi_{\rm out}=10^{4}$.
For later use it is convenient to summarize how the heating rates depend on these model parameters. Near the inner edge, where viscous heating dominates, the vertically integrated viscous flux scales as
$Q_{\rm vis}\propto\dot{m}\,M^{-1}\,(C_{\rm gap}\,a_{0})^{-3},$ (56)
where equation (41) is used. In the irradiation–dominated regime with the $\tau_{3}$ opacity, $Q_{\rm irr}\propto T_{\rm s}^{4}$ with the asymptotic solution of the outer boundary condition gives
$Q_{\rm irr}$ $\propto$ $\alpha_{\rm SS}^{-2/19}\,A^{20/19}\,\dot{m}^{22/19}\,$ (57)
$\times$ $M^{-21/19}\,(C_{\rm gap}a_{0})^{-69/38}\,\xi_{\rm out}^{-69/38},$
where $x_{e}=1$ is adopted for simplicity, and equations (5), (28), and (LABEL:eq:yout3) were used for the derivation. The relative importance of each parameter for the spectral shape is essentially determined by the exponents in the above two scalings.
The strongest dependences in equation (57) arise from the geometric and global scalings, in particular $(C_{\rm gap}a_{0})^{-69/38}$ and subsequently $M^{-21/19}$. Equation (57) also shows that $Q_{\rm irr}(\xi_{\rm out})\propto\xi_{\rm out}^{-69/38}$ for the local irradiation heating rate evaluated near the outer boundary when the other parameters are fixed. This negative exponent reflects that the irradiation flux intercepted per unit area decreases toward larger radii. Nevertheless, $\xi_{\rm out}$ remains important for the observable SED because it sets the maximum emitting radius and therefore the amount of cool reprocessing area included in the radial integration. Increasing $\xi_{\rm out}$ adds cooler annuli that dominate the lowest-frequency emission and broadens the IR bump toward lower frequencies.
The geometric factor $C_{\rm gap}a_{0}$ enters the scalings of both heating terms with a large negative exponent. Because of this scaling, larger values of $C_{\rm gap}a_{0}$ reduce both the viscous heating flux $Q_{\rm vis}$ and the irradiative heating flux $Q_{\rm irr}$ at a given dimensionless radius. In practice, however, $C_{\rm gap}$ is expected to vary only within a narrow range of order unity, whereas the binary separation $a_{0}$ can change by orders of magnitude. For this reason, the variation of $a_{0}$ is much more important than that of $C_{\rm gap}$. We therefore fix $C_{\rm gap}=2$ as a representative value (see equation (3)) and explicitly compare CBD spectra corresponding to $a_{0}=100$ and $a_{0}=1000$ for our four opacity models.
The binary mass $M$ enters the above scalings of both $Q_{\rm vis}$ and $Q_{\rm irr}$ with a moderate exponent. More importantly, $M$ sets $r_{\rm in}$ and determines the global frequency scale of the CBD emission. We therefore explore the dependence on $M$ in Section IV.2, while keeping the other parameters fixed, in order to separate the effect of mass scaling from the opacity-driven changes in the spectral shape.
The parameters $\dot{m}$ and $A$ enter $Q_{\rm irr}$ with moderately large positive exponents, whereas only $\dot{m}$ enters $Q_{\rm vis}$ and it does so linearly. They mostly rescale the overall luminosity of the irradiated disk and only mildly modify the spectral shape in the outer, irradiation dominated region. Similarly, the Shakura–Sunyaev parameter $\alpha_{\rm SS}$ appears in $Q_{\rm irr}$ with a very small exponent and does not enter $Q_{\rm vis}$ explicitly in our formulation, so it has a negligible effect on the spectral shape in this regime. Therefore, for the purpose of illustrating the qualitative differences between the opacity models, we fix $\dot{m}=1$, $A=0.1$, and $\alpha_{\rm SS}=0.1$ in the present work.
Finally, the mass ratio $q$ appears only through $\beta$ and produces only a higher order correction to the irradiation flux. Its influence on the spectral shape of the outer disk is much weaker than that of $\xi_{\rm out}$ or $a_{0}$, and it does not affect $Q_{\rm vis}$ in our prescription. This very weak dependence is consistent with Figure 3(b) of Lee et al. (2024), where the CBD spectra for different mass ratios almost overlap.
Given these scalings, we fix $(q,\alpha_{\rm SS},A,\dot{m},C_{\rm gap})$ to their fiducial values in the rest of this paper. We treat $\xi_{\rm out}$, $M$, and $a_{0}$ as the key parameters that control the CBD spectral shape. In Section IV.2 we present CBD spectra that illustrate the dependence on $M$ and $\xi_{\rm out}$, and we also compare spectra for $a_{0}=1000$ and $a_{0}=100$ to quantify the effect of the binary semi-major axis.
IV Results
@@
IV.1 Circumbinary Disk Temperature Profiles
@@
IV.2 Circumbinary Disk Spectra
@@
$S_{\nu}=\int I_{\nu}\,d\Omega=4\pi\frac{h}{c^{2}}\frac{\cos{\delta}}{D_{\rm L}^{2}}\nu^{3}\int_{r_{\rm in}}^{r_{\rm out}}\,\frac{r}{e^{h\nu/(kT_{\rm s})}-1}\,dr,$
where $\delta$ is the inclination angle of the disk and $D_{\rm L}$ is the luminosity distance. In the following we adopt $\delta=0$ unless otherwise noted.
Figure 4: SEDs of the CBD for the four opacity models listed in Table 1. Panels (a)–(c) correspond to total binary masses of $100\,M_{\odot}$ (fiducial case), $10\,M_{\odot}$, and $1000\,M_{\odot}$, respectively. Each panel plots the normalized luminosity $\nu L_{\nu}/L_{\rm Edd}$ versus frequency on logarithmic axes. Solid curves indicate Model 0 (black), Model 1 (blue), Model 2 (green), and Model 3 (red); the dashed black curve reproduces Model 0 without external irradiation.
Figure 5: Same format as Figure 4 but for different outer disk radii, varying the outer radius $\xi_{\rm out}$ while keeping the binary mass fixed at $M=100\,M_{\odot}$. Panels (a)–(c) adopt $\xi_{\rm out}=10^{4}$ (fiducial case), $10^{3}$, and $10^{5}$, respectively; all other parameters are identical to those in Figure 4.
In what follows, we only vary the model parameters that most strongly affect the CBD spectral shape, namely the binary mass $M$, the dimensionless outer radius $\xi_{\rm out}$, and the dimensionless semi-major axis $a_{0}$, as identified by the scalings of the viscous and irradiation heating rates in Section III.4 (equations (56) and (57)). The remaining parameters $(q,\alpha_{\rm SS},A,\dot{m},C_{\rm gap})$ are kept fixed at their fiducial values, and we examine how the CBD spectra respond to changes in $M$, $\xi_{\rm out}$, and $a_{0}$ for the four opacity prescriptions summarized in Table 1.
@@
Varying the binary mass, $M$, primarily shifts the entire SED along the frequency axis. For a lower-mass system ($M=10\,M_{\odot}$; panel b), both the irradiation-induced IR peak and the viscously heated UV shoulder shift to higher frequencies. In contrast, a higher-mass system ($M=1000\,M_{\odot}$; panel c) produces a cooler and more extended CBD, moving both features to lower frequencies. These mass-dependent shifts are consistent with the scaling relations derived in Section III.4. Because the CBD inner radius scales linearly with $M$ (see equation (3)), the surface temperatures in both the viscously heated inner region and the irradiation-dominated outer region follow the trends implied by equations (56) and (57). In summary, a smaller binary mass gives a more compact and hotter disk whose spectral features move to higher frequencies, whereas a larger binary mass gives a cooler and more extended CBD whose emission peaks at lower frequencies. Thus, the binary mass mainly sets the overall frequency scale of the CBD emission, while the basic spectral shape remains almost unchanged.
At high frequencies ($\nu\gtrsim 10^{15}\,\mathrm{Hz}$), all four models yield nearly identical spectra. In this regime, viscous heating dominates, and the surface temperature is governed by the local energy dissipation rate, making the spectra largely independent of the assumed opacity. The tiny residual offsets in the high-frequency peak arise because frequencies near the peak still receive contributions from radii where irradiation and disk flaring are comparable to viscous heating. At these transition radii, the surface temperatures depend weakly on the opacity model and on the outer boundary condition, shifting the integrated high-frequency luminosity very slightly. At lower frequencies, by contrast, the IR peak does not change monotonically from one model to another. Model 0, which assumes $T_{\mathrm{s}}=T_{\mathrm{c}}$ and neglects all physical opacity sources, gives the brightest IR emission because it sets the surface temperature equal to the hotter midplane temperature and does not include any absorption that would reduce the escaping flux. Among the models that include opacity, Model 3 has the strongest IR flux, followed by Model 2 and then Model 1. This ordering reflects differences in the surface temperature profiles (see Figure 3), which are set by how efficiently each model absorbs and thermalizes the irradiation in the outer part of the CBD.
@@
In Model 3, the inclusion of bound-free absorption greatly increases the Rosseland mean opacity in the temperature range $T\sim 10^{3}$$10^{4}\,\mathrm{K}$, which coincides with the irradiated outer CBD regions where the disk flares (i.e., the aspect ratio increases with radius). This enhances the local trapping of irradiating photons and raises the surface temperature, leading to stronger IR re-emission than in Model 2, where free-free absorption is included together with electron scattering but bound-free absorption is absent. In contrast, electron scattering only in Model 1 does not effectively trap irradiating photons due to the weaker disk flaring, and the surface therefore remains cooler with a lower IR flux. Taken together, these results indicate that observational constraints on CBD opacity sources are most effective in the IR for high-mass binaries and in the optical for low-mass binaries, where the spectral differences among models are most pronounced.
@@
In panel (b) of Figure 4, the spectrum of Model 3 (red curve) clearly exceeds those of the other three models at frequencies $\nu\gtrsim 10^{14.4}\mathrm{Hz}$, spanning the optical and near-UV bands. At lower frequencies ($\nu\lesssim 10^{14.4}\mathrm{Hz}$), Model 0 (black curve), which includes no physical absorption opacity, is the brightest model among the four. Within this panel, the spectra therefore show a clear two-level pattern: Model 0 is brighter at lower frequencies, whereas Model 3 dominates at higher frequencies. Furthermore, when the SEDs of Model 3 are compared across the three mass cases (panels (a), (b), and (c)), the $10M_{\odot}$ case exhibits the highest normalized luminosity.
@@
This behavior is interpreted as follows. In the $10M_{\odot}$ case, the outer surface layers of the disk cool to temperatures below $10^{4}\,\mathrm{K}$, where hydrogen is partially ionized and bound–free opacity is especially strong. High-energy photons emitted by the two minidisks are absorbed in these outer layers, thermalized, and re-emitted at IR wavelengths, which produces a pronounced IR peak. At the same time, the inner disk remains hot enough that its surface temperature does not decrease significantly, and the mid- and high-frequency emission is enhanced. As a result, the SED of Model 3 develops an apparently “triple-peaked” profile, although this shape is best understood as a modification of the basic double-peaked SED structure. The two main peaks are produced by viscous heating in the inner CBD and by reprocessed emission from the irradiated outer disk, and bound–free absorption in Model 3 adds a smaller subpeak in the mid-frequency range by distorting the irradiated component of the spectrum and creating a visible peak between the main IR peak and the UV shoulder. It is therefore most appropriate to regard the SED as fundamentally double-peaked, with an additional mid-frequency peak caused by bound–free opacity effects.